A manufacturer of paper used for packaging requires a minimum strength of 1800 g/cm2. To check on the quality of the paper, a random sample of 13 pieces of paper is selected each hour from the previous hour's production and a strength measurement is recorded for each. The standard deviation σ of the strength measurements, computed by pooling the sum of squares of deviations of many samples, is known to equal 180 g/cm2, and the strength measurements are normally distributed.

a) If the mean of the population of strength measurements is 1850 g/cm2, what is the approximate probability that, for a random sample of n = 13 test pieces of paper,

x < 1800? (Round your answer to four decimal places.)

b) What value would you select for the mean paper strength μ in order that P(x < 1800) be equal to 0.001? (Round your answer to three decimal places.)
Answer #1

(a) Using Central Limit Theorem, standard deviation of the sample,
\sigma _{x}=\frac{\sigma }{\sqrt{n}}=\frac{180}{\sqrt{13}}=49.923

P(x<1800)=P\left ( Z<\frac{1800-1850}{180 /\sqrt{13}} \right )

= P(Z < - 1.0015) = 0.1583 using z table.

(b) Suppose the mean paper strength = μ and P(Z < z1) = 0.001
Using z table, z1 = - 3.090
\left ( \frac{1800-\mu}{180 /\sqrt{13}} \right )=-3.090
or μ = 1954.262